# Aperiodic points for outer billiards

### v3

@@ 3.1. Vassal polygons
Recall that $N$ is even, hence it has the form $N=2m+2$. Let $c_{m}$ be the intersection point of the line $(-\zeta)(-\zeta^{2})$ with the line $1\zeta$. Polygons $\Pi$ and $\Pi^{\dagger}$ are symmetric with respect to the midpoint of the segment $[-\zeta;c_{3}]$. It follows that the line $(-\zeta)(-\zeta^{2})$ contains an edge $[c_{m-1};c_{m}]$ of $\Pi^{\dagger}$. Write $c_{0}$, $\dots$, $c_{m}$ for the vertices of $\Pi^{\dagger}$ symmetric to $1$, $\zeta^{-1}$, $\dots$, $-\zeta$, so that $c_{k}$ is symmetric to $\zeta^{-k}$. Both $\Pi$ and $\Pi^{\dagger}$ are symmetric with respect to the reflection axis of strip 0. Hence the line $1\zeta^{-1}$ contains the edge $[c_{0};c_{1}]$ of $\Pi^{\dagger}$.
@@
Note that any diameter line of $\Pi$ is a symmetry axis of $X$, and that $f_{\Pi}$ takes $A_{0}$ to the symmertric polygon with respect to this axis. In particular, $f_{\Pi}(A_{0})\subset X$, and similarly for all pieces $A_{k}:=R^{k}A_{0}$ for $k=0$, $\dots$ $N-1$. The restriction of $f_{\Pi}$ to $X\setminus\Pi$ is therefore a piecewise rotation of $X$ with pieces $A_{k}$, cf. [Ruk22, Lemma 20]. Denote it by $f_{X}$.
@@ 3.2. Computation of $\mathrm{Inv}_{L}(g_{X})$
All tensor products we consider are over $\mathbb{Q}$; for this reason, we will from now on omit $\mathbb{Q}$ from the notation. For our subsequent computation of the dynamic Hadwiger invariant for $g_{X}$, we will need to know the lengths of all edges of $A_{0}$. Edges $[c_{1};c_{2}]$, $\dots$, $[c_{m-1};c_{m}]$ all have the same length as an edge of $\Pi$. Namely, their length is $\lambda_{0}=2\sin\frac{\theta}{2}$. Side edges $[1,c_{1}]$ and $[1;c_{m}]$ of $A_{0}$ have the same length $\lambda^{\prime}$, which we now want to compute. Consider the triangle with vertices $1$, $-1$, $c_{0}$; this is an isosceles triangle with base length 2 and angles $\theta$, $\frac{\pi-\theta}{2}$, $\frac{\pi-\theta}{2}$. It follows that the side edge of this triangle has length
@@
A dynamical Hadwiger valuation of $g_{X}$ is associated with some choice of a line $L$. Choose the line $L$ through the vertices $1=\zeta^{0}$ and $\zeta^{1}$ of $\Pi$, and define the positive side of $L$ as the one not containing $\Pi$. Each edge $e$ of $A_{0}$ gives rise to two $L$-adapted pairs (see Section 2.5) $(A_{k},E)$ and $(MA_{k},ME)$, where $E=R^{k}e$ is the corresponding to $e$ edge of $A_{k}$ parallel to $L$. Call the pairs $(A_{k},E)$ and $(MA_{k},ME)$ the $L$-adapted pairs corresponding to $e$. Among these two pairs, exactly one is positive, and the other one is negative. For example, $(A_{0},[c_{m};1])$ is a negative pair, hence, the corresponding to the edge $[c_{m};1]$ positive pair is $(A_{m+1},M[c_{m},1])$. Positive pairs are described in the next lemma.
@@
By Theorem 2.29, a periodic polytope exchange transformation of a bounded polytope must have zero invariant $\mathrm{Inv}_{L}$. It now follows from Corollary 3.7 that $g_{X}$ is not periodic. Hence $f_{X}=M\circ g_{X}$ is not periodic either, since $M$ has period 2 and commutes with $f_{X}$ and $g_{X}$. We have thus proved the following theorem.
@@ 4. Boundary dynamics of piecewise rotations
Section 4 describes alternative methods for proving similar results to those of Section 3. While it can be useful for future research, this section is not directly involved in the proof of Theorem 1.1 and, for this reason, can be omitted. Given a finite piecewise rotation $f$ of a bounded polygon $P$, we will introduce an associated interval exchange transformation. Computing the SAF invariant of the latter transformation helps determining aperiodicity of the initial map $f$. Let $P_{1}$, $\dots$, $P_{m}$ be the components of $\mathrm{dom}(f)$. Define $\partial f$ as the disjoint union of the boundaries of all $P_{i}$. More precisely, elements of $\partial f$ are pairs of the form $(z,P_{i})$, where $z\in\partial P_{i}$. Then $\partial f$ is a finite graph that is a union of $m$ cycles. It has a natural orientation induced from the orientation of $\mathbb{R}^{2}$: every component of $\partial f$ corresponding to the boundary of some piece $P_{i}$ is oriented as the boundary, i.e., counterclockwise.
Note that $\partial f$ projects naturally to $\overline{P}$; this projection is two-to-one except the points projecting to $\partial P$ and except finitely many vertices of $\partial f$. Every point of $\partial P$ that is not a vertex of a component of $\mathrm{dom}(f)$ has only one preimage in $\partial f$, and vertices of pieces $P_{i}$ may have more than two preimages.
@@ 4.1. Boundary quasi-commutators
Recall that, when dealing with interval exchange transformations, one ignores any effect happening only on a finite subset of points. In particular, it makes sense to consider the group $\mathrm{IET}(\partial f)$ of all interval exchange transformations of $\partial f$, even though $\partial f$ is not an interval (but one can make it into an interval by cutting and pasting at finitely many points). Note that $f$ defines a natural map $f^{*}:\partial f\to\partial f^{-1}$. More precisely, $f^{*}$ sends a point $(z,i)\in\partial f$ to the point $(\overline{f}_{i}(z),f(P_{i}))$, where $\overline{f}_{i}$ is the continuous extension of $f$ to the closure of $P_{i}$. A disadvantage of $f^{*}$ is that this map cannot be iterated. However, below, we will define an iterable map based on $f^{*}$.
@@
Let $\partial_{in}f$ be the inner boundary of $f$, formed by (the copies of) the edges of $P_{i}$ not on the boundary of $P$. Call the remaining part of $\partial f$ the outer boundary and denote it by $\partial_{out}f$. Both $\partial_{in}f$ and $\partial_{out}f$ are viewed as subgraphs of $\partial f$. There is a natural involution $w:\partial_{in}f\to\partial_{in}f$ that takes $(z,i)$ to $(z,j)$ whenever $z$ is a common boundary point of $P_{i}$ and $P_{j}$. Informally speaking, $w$ changes sides. Note that $w$ is undefined at vertices of $P_{i}$ that lie inside $P$, however, this is immaterial for the definition of $v$ as a piecewise isometry of $\partial_{in}f$. Clearly, $w\circ w=id_{\partial_{in}f}$, except at finitely many points. Also, observe that $w$ reverses the orientation of $\partial_{in}$. Extend $w$ to all of $\partial f$ as an involutive orientation reversing piecewise isometry. This is always possible: one just needs to define $w$ on $\partial P$ (except at finitely many points, $\partial P$ identifies with a subset of $\partial f$), for example, realizing $\partial P$ as a circle and letting $w$ act on this circle as a reflection with respect to some axis. A map $w$ thus extended is called a boundary involution of $f$. By definition, it reverses the orientation and satisfies $w\circ w=id_{\partial f}$, where, as always, we ignore finite sets of points.
Write $\mathrm{IETF}(\partial f)$ for the group of all Euclidean scissors automorphisms of the graph $\partial f$, not necessarily orientation preserving. A boundary involution $w$ introduced above is an element of $\mathrm{IETF}(\partial f)$. Let $w_{-1}$ be the boundary involution for $f^{-1}$, where the action of $w_{-1}$ is chosen to coincide with $w$ on $\partial P$. With this proviso, $w_{-1}$ is uniquely determined by $w$. Now, we can define a non-trivially iterable interval exchange transformation associated with $f^{*}$.
@@
Let $f$ be a $\mathsf{G}$-scissors congruence of a compact polygon $P$, where $\mathsf{G}$ is the group of all orientation preserving Euclidean isometries. If $f$ is periodic, then any associated boundary quasi-commutator has zero SAF invariant.
@@ 4.2. Induced outer billiard map
Let $\pi:\mathrm{dom}(f_{\Pi})\to V_{0}$ be the map sending $V_{k}:=R^{k}V_{0}$ to $V_{0}$ by $R^{-k}$, for every $k=0$, $\dots$, $N-1$. Then the induced outer billiard map $f_{\mathrm{ind}}$ is defined as $\pi\circ f_{\Pi}$.
@@
We start with describing level one pieces of $f_{\mathrm{ind}}$. The number $N$ of edges of $\Pi$ is not assumed to be even. Set $m$ to be $[N/2]+1$, where $[N/2]$ is the integer floor of $N/2$, which is consistent with our earlier use of $m$ in the even case. Let $v_{0}$ stand for the reflection (=axial symmetry) of $V_{0}$ through the vertices axis passing through point 1; the axis of $v_{0}$ is shown as a dashed vertical line in Fig. 4.
@@
First, we make several observations. The intersection $V_{0}\cap S_{-k}V_{-k}$ is nonempty if and only if $k<\frac{N}{2}+1$, that is, if $k\leqslant m+1$. It follows that $V_{0}\cap S_{-k}V_{-k}$, for $k=1$, $\dots$, $m+1$, tile the sector $V_{0}$. Another observation is that, on $V_{0}$, the map $f_{\Pi}=S_{0}$ acts as $u_{0}\circ v_{0}$, where $u_{0}$ is the reflection in the real line, that is, the complex conjugation. Observe, finally, that $u_{0}$ takes the sector $S_{-k}V_{-k}$ to the sector $V_{k}$. Indeed, since $V_{-k}$ has vertex $\zeta^{-k}$, the sector $u_{0}S_{-k}V_{-k}$ has vertex $\zeta^{k}=u_{0}(\zeta^{-k})$. Now, it only remains to figure out which of the two sectors with vertex $\zeta^{k}$ is represented by $u_{0}V_{-k}$ and which by $u_{0}S_{-k}V_{-k}$; this is straighforward by noticing that $V_{-k}$ and $u_{0}S_{-k}V_{-k}$ are oriented clockwise with respect to the origin.
@@ 4.3. Geometry of the pieces $Q_{k}$
Below, we compute geometric parameters of the pieces $Q_{k}$ such as the angles and the sidelengths, i.e., lengths of the edges. Observe that the pieces $Q_{1}$ through $Q_{m/2}$ form a triangle $U$ that is invariant under $f_{\mathrm{ind}}$ in the sense that, for every piece $Q_{k}\subset U$, we also have $f_{\mathrm{ind}}(Q_{k})\subset U$. The triangle $U$ is isosceles with horizontal base; the existence of such a triangle depends on our assumption that $m$ is even. In particular, the restriction $f_{U}$ of $f_{\mathrm{ind}}$ to $U$ is a piecewise rotation of $U$.
Denote the vertices of the level 1 pieces of $f_{U}$ that lie on the left edge of $U$ by $a_{1}$, $\dots$, $a_{m/2}$; the numeration is downward. Similarly, let $b_{1}$, $\dots$, $b_{m/2}$ be the vertices on the right edge of $U$; see Fig. 5. In particular, $Q_{1}$ is the triangle with vertices $\zeta^{0}$, $a_{1}$, $b_{1}$, and each $Q_{k}$ with $k>1$ is a quadrilateral with vertices $a_{k}$, $b_{k}$, $b_{k-1}$, $a_{k-1}$, where the vertices are always listed in a counterclockwise order. Note also that $v$ acts on $Q_{k}$ by fixing the vertices $a_{k}$, $b_{k-1}$ while interchanging $a_{k-1}$, $b_{k}$.
@@ 4.4. Computing boundary SAF invariants
A boundary SAF invariant of a piecewise rotation $f$ is defined as the SAF invariant of a boundary quasi-commutator of $f$. Computing boundary SAF invariants for $f$ helps establishing aperiodicity. Below, we give a specific example of such a computation. Namely, consider the induced billiard map $f_{\mathrm{ind}}$ in the case when $N=2m+2$ is even but not divisible by 4; in this case, $m$ is also even. As follows from the above, $U$ is an invariant triangle of $f_{\mathrm{ind}}$, and $f_{U}$ stands for the restriction of $f_{\mathrm{ind}}$ to $U$. We will compute a boundary SAF invariant of $f_{U}$. First, one needs to define the boundary involution on the external boundary of $U$. A standard choice is the axial symmety of $U$, that is, the reflection $v_{0}$ of $U$ in the vertical axis passing through point $1$. Both the boundary involution and the boundary quasi-commutator are now well defined for $f_{U}$.
@@
a property that plays a definite role in what follows. Let $u^{*}$ and $v^{*}$ be the induced maps of $\partial f$; the former sending $\partial f$ to $\partial f^{-1}$, and the latter — to $\partial f$ itself. Also, write $f^{*}$ instead of $f^{*}_{\mathrm{ind}}$ to lighten the notation. Since $f_{\mathrm{ind}}=u\circ v$ by Lemma 4.7, one also has $f^{*}=u^{*}\circ v^{*}$.
@@
Let $F$ be a partially defined map of a set $D$, that is, $F:\mathrm{dom}(F)\to D$, where $\mathrm{dom}(F)\subset D$. For any subset $A\subset D$ and any point $x\in A$, the first return time $n_{x}$ to $A$ is defined as the smallest positive integer $n$ with the property $F^{n}(x)\in A$ or infinity if such $n>0$ does not exist. By definition, the domain of the first return map $\mathcal{R}_{A}F$ consists of all points $x\in A$ such that $n_{x}<\infty$. Every $x\in\mathrm{dom}(\mathcal{R}_{A}F)$ is mapped to the point $\mathcal{R}_{A}F(x)=F^{n_{x}}(x)$.
In our case, $D=\partial_{out}f_{U}$, and $F=v^{*}\circ w$. The first return map $\mathcal{R}_{out}$ can therefore be viewed as an interval exchange transformation of $\partial U$, upon the identification between $\partial U$ and $\partial_{out}f_{U}$. We can explicitly describe the map $\mathcal{R}_{out}$.
@@
Define a map $\rho\in\mathrm{IETF}(\partial U)$ as follows. The domain of $\rho$ will be the complement in $\partial U$ to finitely many points; namely, all vertices of $U$ and all vertices of all $Q_{i}$, the components of $\mathrm{dom}(f_{U})$. Now, take a point $z\in\mathrm{dom}(\rho)$; then $z$ lies on the boundary of a (necessarily unique) component $Q_{i}$. There are two cases: either $\overline{v}_{i}(z)$ lies in $\partial U$, or not (the latter case means that $\overline{v}_{i}(z)$ is a common boundary point of two pieces of $f_{U}$). Here, $\overline{v}_{i}$ is the continuous extension of $v_{i}$ to $\overline{Q}_{i}$, the closure of $Q_{i}$. In the former case, set $\rho(z):=\overline{v}_{i}(z)$. Consider the latter case: there is a unique $Q_{j}\neq Q_{i}$ such that $\overline{v}_{i}(z)\in\partial Q_{j}$. Set $\rho(z)=\overline{v}_{j}\circ\overline{v}_{i}(z)$; this point necessarily lies in $\partial U$. In terms of the dynamics on the graph $\partial f_{U}$, the map corresponding to $\rho$ can be written as $v^{*}\circ w\circ v^{*}$ in the latter case and as $v^{*}$ in the former case.
It is not hard to describe the map $\rho$ in our specific situation: $\rho$ acts on $\partial U$ as a reflection with fixed point $a_{m/2}$; the other fixed point is the midpoint $b_{0}$ of the segment $[1;b_{1}]$, that is, the point where the symmetry axis of $P_{1}$ crosses the edge $[1;b_{1}]$. Recall that this is an orientation reversing isometry of $\partial U$ (with respect to the length metric) which is conjugate to a Euclidean reflection under any isometric homeomorphism between $\partial U$ and a circle. Explicitly, $\rho$ interchanges the following pairs of straight line segments:
@@
Recall that $\overline{u}$, the continuous extension of $u$ to $\overline{U}$, also acts as a reflection of $\partial U$; its fixed points are $1$ and the midpoint of $[a_{m/2};b_{m/2}]$. As a composition of two reflections, $\mathcal{R}_{out}=\rho\circ u^{*}$ is necessarily a rotation by some angle; it remains only to compute the angle $\alpha$ of this rotation. Note that the point $b_{1}$ is mapped to the point $1$; hence
@@
It remains to find a relationship between the SAF invariants of the interval exchange maps $v^{*}\circ w$ and $\mathcal{R}_{out}$. Recall that an interval swap map is an interval exchange transformation $s$ of some graph that is different from the identity only on a disjoint union of two intervals, say, $I_{1}$ and $I_{2}$, and such that
  • on each of the two intervals, $s$ acts as a translation;
  • it takes $I_{1}$ onto $I_{2}$ and $I_{2}$ onto $I_{1}$.
By definition of the SAF invariant, it is immediate that $\mathrm{Inv}(s)=0$ for every interval swap map.

### v4

@@ 3.1. Vassal polygons
Recall that $N$ is even, hence it has the form $N=2m+2$. Let $c_{m}$ be the intersection point of the line $(-\zeta)(-\zeta^{2})$ with the line $1\zeta$. Polygons $\Pi$ and $\Pi^{\dagger}$ are symmetric with respect to the midpoint of the segment $[-\zeta;c_{m}]$. It follows that the line $(-\zeta)(-\zeta^{2})$ contains an edge $[c_{m-1};c_{m}]$ of $\Pi^{\dagger}$. Write $c_{0}$, $\dots$, $c_{m}$ for the vertices of $\Pi^{\dagger}$ symmetric to $1$, $\zeta^{-1}$, $\dots$, $\zeta^{-m}=-\zeta$, so that $c_{k}$ is symmetric to $\zeta^{-k}$. Both $\Pi$ and $\Pi^{\dagger}$ are symmetric with respect to the reflection axis of strip 0. Hence the line $1\zeta^{-1}$ contains the edge $[c_{0};c_{1}]$ of $\Pi^{\dagger}$.
@@
Note that any diameter line of $\Pi$ is a symmetry axis of $X$, and that $f_{\Pi}$ takes $A_{0}$ to the symmetric polygon with respect to the diameter through $1$ and $-1$. In particular, $f_{\Pi}(A_{0})\subset X$, and similarly for all pieces $A_{k}:=R^{k}A_{0}$ for $k=0$, $\dots$ $N-1$. The restriction of $f_{\Pi}$ to $X\setminus\Pi$ is therefore a piecewise rotation of $X$ with pieces $A_{k}$, cf. [Ruk22, Lemma 20]. Denote it by $f_{X}$.
@@ 3.2. Computation of $\mathrm{Inv}_{L}(g_{X})$
All tensor products we consider are over $\mathbb{Q}$; for this reason, we will from now on omit $\mathbb{Q}$ from the notation. For our subsequent computation of the dynamic Hadwiger invariant for $g_{X}$, we will need to know the lengths of all edges of $A_{0}$. Edges $[c_{1};c_{2}]$, $\dots$, $[c_{m-1};c_{m}]$ all have the same length as an edge of $\Pi$. Namely, their length is $\lambda_{0}=2\sin\frac{\theta}{2}$. Side edges $[1,c_{1}]$ and $[1;c_{m}]$ of $A_{0}$ have the same length $\lambda^{\prime}$, which we will now compute. Consider the triangle with vertices $1$, $-1$, $c_{0}$; this is an isosceles triangle with base length 2 and angles $\theta$, $\frac{\pi-\theta}{2}$, $\frac{\pi-\theta}{2}$. It follows that the side edge of this triangle has length
@@
A dynamic Hadwiger invariant of $g_{X}$ is associated with some choice of a line $L$. Choose the line $L$ through the vertices $1=\zeta^{0}$ and $\zeta^{1}$ of $\Pi$, and define the positive side of $L$ as the one not containing $\Pi$. Each edge $e$ of $A_{0}$ gives rise to two $L$-adapted pairs (see Section 2.6) $(A_{k},E)$ and $(M(A_{k}),M(E))$, where $E=R^{k}e$ is the corresponding to $e$ edge of $A_{k}$ parallel to $L$, and $M$ is the map $z\to-z$. Call the pairs $(A_{k},E)$ and $(MA_{k},ME)$ the $L$-adapted pairs corresponding to $e$. Among these two pairs, exactly one is positive, and the other one is negative. For example, $(A_{0},[c_{m};1])$ is a negative pair, hence, the corresponding to the edge $[c_{m};1]$ positive pair is $(A_{m+1},M[c_{m},1])$. Positive pairs are described in the next lemma.
@@
By Theorem 2.29, a periodic polygon exchange transformation of a bounded polygon must have zero invariant $\mathrm{Inv}_{L}$. It now follows from Corollary 3.7 that $g_{X}$ is not periodic. Hence $f_{X}=M\circ g_{X}$ is not periodic either, since $M$ has period 2 and commutes with $f_{X}$ and $g_{X}$. We have thus proved the following theorem.
@@ 4. Boundary dynamics of piecewise rotations
Section 4 describes alternative methods for proving similar results to those of Section 3. While it can be useful for future research, this section is not directly involved in the proof of Theorem 1.1 and, for this reason, can be omitted. Given an orientation preserving scissors automorphism $f$ of a bounded polygon $P$, we will introduce an associated interval exchange transformation. Computing the SAF invariant of the latter transformation helps determining aperiodicity of the initial map $f$. Let $P_{1}$, $\dots$, $P_{m}$ be the components of $\mathrm{dom}(f)$. Define $\partial f$ as the disjoint union of the boundaries of all $P_{i}$. More precisely, elements of $\partial f$ are pairs of the form $(z,P_{i})$, where $z\in\partial P_{i}$. Then $\partial f$ is a finite graph that is a union of $m$ cycles. It has a natural orientation induced from the orientation of $\mathbb{R}^{2}$: every component of $\partial f$ corresponding to the boundary of some piece $P_{i}$ is oriented as the boundary, i.e., counterclockwise.
Note that $\partial f$ projects naturally onto $\overline{P}\setminus\mathrm{dom}(f)$; this projection is two-to-one except the points projecting to $\partial P$ and except finitely many vertices of $\partial f$. Every point of $\partial P$ that is not a vertex of a component of $\mathrm{dom}(f)$ has only one preimage in $\partial f$, and vertices of pieces $P_{i}$ may have more than two preimages.
@@ 4.1. Boundary quasi-commutators
Recall that, when dealing with interval exchange transformations, one ignores any effect happening only on a finite subset of points. In particular, it makes sense to consider the group $\mathrm{IET}(\partial f)$ of all interval exchange transformations of $\partial f$, even though $\partial f$ is not an interval (but one can make it into an interval by cutting and pasting at finitely many points). Note that $f$ defines a natural map $f^{*}:\partial f\to\partial f^{-1}$. More precisely, $f^{*}$ sends a point $(z,i)\in\partial f$ to the point $(\overline{f}_{i}(z),f(P_{i}))$, where $\overline{f}_{i}$ is the continuous extension of $f|_{P_{i}}$ to the closure of $P_{i}$. A disadvantage of $f^{*}$ is that this map cannot be iterated. However, below, we will define an iterable map based on $f^{*}$.
@@
Let $\partial_{in}f$ be the inner boundary of $f$, formed by (the copies of) the edges of $P_{i}$ not on the boundary of $P$. Call the remaining part of $\partial f$ the outer boundary and denote it by $\partial_{out}f$. Both $\partial_{in}f$ and $\partial_{out}f$ are viewed as subgraphs of $\partial f$. There is a natural involution $w:\partial_{in}f\to\partial_{in}f$ that takes $(z,i)$ to $(z,j)$ whenever $z$ is a common boundary point of $P_{i}$ and $P_{j}$. Informally speaking, $w$ changes sides. Note that $w$ is undefined at vertices of $P_{i}$ that lie inside $P$, however, this is immaterial for the definition of $w$ as a piecewise isometry of $\partial_{in}f$. Clearly, $w\circ w=id_{\partial_{in}f}$, except at finitely many points. Also, observe that $w$ reverses the orientation of $\partial_{in}f$. Extend $w$ to all of $\partial f$ as an involutive orientation reversing piecewise isometry. This is always possible: one just needs to define $w$ on $\partial P$ (except at finitely many points, $\partial P$ identifies with $\partial_{out}f$), for example, realizing $\partial P$ as a circle and letting $w$ act on this circle as a reflection with respect to some axis. A map $w$ thus extended is called a boundary involution of $f$. By definition, it reverses the orientation and satisfies $w\circ w=id_{\partial f}$, where, as always, we ignore finite sets of points.
Write $\mathrm{IETF}(\partial f)$ for the group of all Euclidean scissors automorphisms of the graph $\partial f$, not necessarily orientation preserving; these automorphisms are also known as interval exchange transformations with flips, hence the abbreviation IETF. A boundary involution $w$ introduced above is an element of $\mathrm{IETF}(\partial f)$. Let $w_{-1}$ be the boundary involution for $f^{-1}$, where the action of $w_{-1}$ is chosen to coincide with $w$ on $\partial P$. With this proviso, $w_{-1}$ is uniquely determined by $w$. Now, we can define a non-trivially iterable interval exchange transformation associated with $f^{*}$.
@@
Let $f$ be a $\mathsf{G}$-scissors automorphism of a compact polygon $P$, where $\mathsf{G}$ is the group of all orientation preserving Euclidean isometries. If $f$ is periodic, then any associated boundary quasi-commutator has zero SAF invariant.
@@ 4.2. Induced outer billiard map
Let $\pi:\mathrm{dom}(f_{\Pi})\to V_{0}$ be the map sending $V_{k}:=R^{k}V_{0}$ to $V_{0}$ by $R^{-k}$, for every $k=0$, $\dots$, $N-1$. Then the induced outer billiard map $f_{\mathrm{ind}}:V_{0}\cap\mathrm{dom}(f_{\Pi}^{2})\to V_{0}$ is defined as $\pi\circ f_{\Pi}$.
@@
We start with describing level one pieces of $f_{\mathrm{ind}}$. The number $N$ of edges of $\Pi$ is not assumed to be even. Set $m$ to be $[\frac{N+1}{2}]-1$, where $[\frac{N+1}{2}]$ is the integer floor of $\frac{N+1}{2}$, which is consistent with our earlier use of $m$ in the even case. Let $v_{0}$ stand for the reflection (=axial symmetry) of $V_{0}$ through the vertices axis passing through point 1; the axis of $v_{0}$ is shown as a dashed vertical line in Fig. 4.
@@
First, we make several observations. The intersection $V_{0}\cap S_{-k}V_{-k}$ is nonempty if and only if $k<\frac{N+1}{2}$, that is, if $k\leqslant m+1$. It follows that $V_{0}\cap S_{-k}V_{-k}$, for $k=1$, $\dots$, $m+1$, tile the sector $V_{0}$. Another observation is that, on $V_{0}$, the map $f_{\Pi}=S_{0}$ acts as $u_{0}\circ v_{0}$, where $u_{0}$ is the reflection in the real line, that is, the complex conjugation. Observe, finally, that $u_{0}$ takes the sector $S_{-k}V_{-k}$ to the sector $V_{k}$. Indeed, since $V_{-k}$ has vertex $\zeta^{-k}$, the sector $u_{0}S_{-k}V_{-k}$ has vertex $\zeta^{k}=u_{0}(\zeta^{-k})$. Now, it only remains to figure out which of the two sectors with vertex $\zeta^{k}$ bounded by support lines of $\Pi$ is represented by $u_{0}V_{-k}$ and which by $u_{0}S_{-k}V_{-k}$; this is straighforward by noticing that $V_{-k}$ and $u_{0}S_{-k}V_{-k}$ are oriented clockwise with respect to the origin.
@@ 4.3. Geometry of the pieces $Q_{k}$
In Section 4.3, we assume that $N=2m+2$ is twice odd, and that $m$ is even, that is, $N\equiv 2$ modulo 4. Below, we compute geometric parameters of the pieces $Q_{k}$ such as the angles and the sidelengths, i.e., lengths of the edges. Observe that the pieces $Q_{1}$ through $Q_{m/2}$ form a triangle $U$ that is invariant under $f_{\mathrm{ind}}$ in the sense that, for every piece $Q_{k}\subset U$, we also have $f_{\mathrm{ind}}(Q_{k})\subset U$. The triangle $U$ is isosceles with horizontal base; the existence of such a triangle depends on our assumption that $m$ is even. In particular, the restriction $f_{U}$ of $f_{\mathrm{ind}}$ to $U$ is an orientation preserving scissors automorphism of $U$.
Denote the vertices of the level 1 pieces of $f_{U}$ that lie on the left edge of $U$ by $a_{1}$, $\dots$, $a_{m/2}$; the numeration is downward. Similarly, let $b_{1}$, $\dots$, $b_{m/2}$ be the vertices on the right edge of $U$; see Fig. 5. In particular, $Q_{1}$ is the triangle with vertices $\zeta^{0}$, $a_{1}$, $b_{1}$, and each $Q_{k}$ with $k>1$ is a quadrilateral with vertices $a_{k}$, $b_{k}$, $b_{k-1}$, $a_{k-1}$, where the vertices are always listed in a counterclockwise order. Note also that the reflection $v_{k}$ acts on $Q_{k}$ by fixing the vertices $a_{k}$, $b_{k-1}$ while interchanging $a_{k-1}$, $b_{k}$.
@@ 4.4. Computing boundary SAF invariants
A boundary SAF invariant of an orientation preserving scissors automorphism $f$ is defined as the SAF invariant of a boundary quasi-commutator of $f$. Computing boundary SAF invariants for $f$ helps establishing aperiodicity. Below, we give a specific example of such a computation. Namely, consider the induced billiard map $f_{\mathrm{ind}}$ in the case when $N=2m+2$ is even but not divisible by 4; in this case, $m$ is also even. As follows from the above, $U$ is an invariant triangle of $f_{\mathrm{ind}}$, and $f_{U}$ stands for the restriction of $f_{\mathrm{ind}}$ to $U$. We will compute a boundary SAF invariant of $f_{U}$. First, one needs to define the boundary involution on the external boundary of $U$. A standard choice is the axial symmety of $U$, that is, the reflection $v_{0}$ of $U$ in the vertical axis passing through point $1$. Both the boundary involution and the boundary quasi-commutator are now well defined for $f_{U}$.
@@
a property that plays a definite role in what follows. Let $u^{*}$ and $v^{*}$ be the induced by $u$ and $v$, respectively, maps of $\partial f$; the former sending $\partial f$ to $\partial f^{-1}$, and the latter — to $\partial f$ itself. Also, write $f^{*}$ instead of $f^{*}_{\mathrm{ind}}$ to lighten the notation. Since $f_{\mathrm{ind}}=u\circ v$ by Lemma 4.7, one also has $f^{*}=u^{*}\circ v^{*}$.
@@
Let $F$ be a partially defined map of a set $D$ to itself, that is, $F:\mathrm{dom}(F)\to D$, where $\mathrm{dom}(F)\subset D$. For any subset $A\subset D$ and any point $x\in A$, the first return time $n_{x}$ to $A$ is defined as the smallest positive integer $n$ with the property $F^{n}(x)\in A$ or infinity if such $n>0$ does not exist. By definition, the domain of the first return map $\mathcal{R}_{A}F$ consists of all points $x\in A$ such that $n_{x}<\infty$. Every $x\in\mathrm{dom}(\mathcal{R}_{A}F)$ is mapped to the point $\mathcal{R}_{A}F(x)=F^{n_{x}}(x)$.
In our case, $D=\partial f_{U}$, $A=\partial_{out}f_{U}$, and $F=v^{*}\circ w$. The first return map $\mathcal{R}_{out}:=\mathcal{R}_{A}F$ can therefore be viewed as an interval exchange transformation of $\partial U$, upon the identification between $\partial U$ and $\partial_{out}f_{U}$. We can explicitly describe the map $\mathcal{R}_{out}$.
@@
Recall that $\overline{v}_{i}$ is the continuous extension of $v_{i}$ to $\overline{Q}_{i}$, the closure of $Q_{i}$. Similarly, write $\overline{u}$ for the continuous extension of $u$ to $\overline{U}$.
Define a map $\rho\in\mathrm{IETF}(\partial U)$ as follows. The domain of $\rho$ will be the complement in $\partial U$ of finitely many points; namely, all vertices of $U$ and all vertices of all $Q_{i}$, the components of $\mathrm{dom}(f_{U})$. Now, take a point $z\in\mathrm{dom}(\rho)$; then $z$ lies on the boundary of a (necessarily unique) component $Q_{i}$. There are two cases: either $\overline{v}_{i}(z)$ lies in $\partial U$, or not (the latter case means that $\overline{v}_{i}(z)$ is a common boundary point of two pieces of $f_{U}$). In the former case, set $\rho(z):=\overline{v}_{i}(z)$. Consider the latter case: there is a unique $Q_{j}\neq Q_{i}$ such that $\overline{v}_{i}(z)\in\partial Q_{j}$. Set $\rho(z)=\overline{v}_{j}\circ\overline{v}_{i}(z)$; this point necessarily lies in $\partial U$. In terms of the dynamics on the graph $\partial f_{U}$, the map corresponding to $\rho$ can be written as $v^{*}\circ w\circ v^{*}$ in the latter case and as $v^{*}$ in the former case.
It is not hard to describe the map $\rho$ in our specific situation: $\rho$ acts on $\partial U$ as a reflection with fixed point $a_{m/2}$; the other fixed point is the midpoint $b_{0}$ of the segment $[1;b_{1}]$, that is, the point where the symmetry axis of $Q_{1}$ crosses the edge $[1;b_{1}]$. Recall that this is an orientation reversing isometry of $\partial U$ (with respect to the length metric) which is conjugate to a Euclidean reflection under any isometric homeomorphism between $\partial U$ and a circle. Explicitly, $\rho$ interchanges the following pairs of straight line segments:
@@
Recall that $\overline{u}$ also acts as a reflection of $\partial U$; its fixed points are $1$ and the midpoint of $[a_{m/2};b_{m/2}]$. As a composition of two reflections, $\mathcal{R}_{out}=\rho\circ u^{*}$ is necessarily a rotation by some angle; it remains only to compute the angle $\alpha$ of this rotation. Note that the point $b_{1}$ is mapped to the point $1$; hence
@@
It remains to find a relationship between the SAF invariants of the interval exchange maps $v^{*}\circ w$ and $\mathcal{R}_{out}$. Recall that an interval swap map was defined in Section 2.3, see Example 2.12. By definition of the SAF invariant, it is immediate that $\mathrm{Inv}(s)=0$ for every interval swap map.